OpenStudy (anonymous):

What number must be added to the polynomial to complete the square? If necessary, use the slash mark (/)as a fraction bar and leave your answer in the form of an improper fraction. x^2-9x

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

ok for this one it would be -9=3^2=9?

OpenStudy (anonymous):

Well, you would take HALF of nine. 3 + 3 isn't nine.

OpenStudy (anonymous):

Instead, you would do (9/2)^2

OpenStudy (anonymous):

6^2=36

OpenStudy (anonymous):

\[(\frac{ 9 }{ 2 })^{2}\]

OpenStudy (anonymous):

final answer 9/2

OpenStudy (anonymous):

So 81/4, as a fraction.

OpenStudy (anonymous):

81/4 huh answer

OpenStudy (anonymous):

Yes, x^2 - 9x + 81/4